Four cards are drawn without replacement from pack of cards . What is the probability that (a) they all are of different suit .
(b) they all are aces.

Answer by stanbon(75887)

Four cards are drawn without replacement from pack of cards .
What is the probability that (
a) they all are of different suit.
# of ways to succeed: 13*13*13*13 = 28561
# of random results: 52C4 = 270725
P(4 of different suit) = 13^4/52C4 = 0.1055
------------------------------------
b) they all are aces.
# of ways to succeed: 1
# of random results 52C4
---
P(4 aces) = 1/52C4 = 0.000003694
====================================
Cheers,
Stan H.
==============